good day
sorry could not find Calculus in the list. It's differentiation.

Determine the gradient of the following
g(x) = (2x² / 3) -²
the answer is -9/2 x^-3

Answer by rothauserc(4718)

the derivative of g(x) = (2x² / 3) -² is calculated
g(x) = (2^-2 * x^-4) / 3^-2
g(x) = 9 / 4x^4
g(x) = 9 * x^-4 / 4
g'(x) = (-4 * 9 * x^-5) / 4
g'(x) = -9 / x^5
*************************************************************************
gradient can be the slope or in vector analysis it is the vector operator
also the gradient function is dy/dx or g'(x) for your problem
*************************************************************************
I suspect you want the gradient function